OpenStudy (anonymous):

Is (-6,6) a solution of 5x-7y=2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Have you looked at my questions agdgdgdgwngo?

OpenStudy (anonymous):

Put -6 for x and 6 for y in the equation. Then, 5*(-6)-7*6=-30-42=-72 So it is not the solution for the equation